OpenStudy (anonymous):

1. Using the Chain Rule, fi nd dw/dt when: w = 2ye^(x) - ln z; x = ln (t^(2) + 1) ; y = tan t; z = e^(t)

OpenStudy (mathstudent55):

\(w = 2ye^x - \ln z\) \(w = 2(\tan t) e^{ ln (t^2 + 1)} - \ln (e^t)\)

OpenStudy (mathstudent55):

In general, what is \(e^{\ln x} \) equal to?

OpenStudy (anonymous):

x

OpenStudy (mathstudent55):

Also, what is \( \ln e^x\) equal to?

OpenStudy (anonymous):

1

OpenStudy (mathstudent55):

No, it's also x.

OpenStudy (mathstudent55):

Using those two substitutions will make your expression much simpler. Then you can differentitate it.

OpenStudy (mathstudent55):

\(w = 2(\tan t) e^{ ln (t^2 + 1)} - \ln (e^t)\) First, substitute \(t^2 + 1 \) for \(e^{ ln (t^2 + 1)}\) and \(t \) for \(\ln (e^t)\).

OpenStudy (anonymous):

w = (2t^2+2)(tant)-t

OpenStudy (mathstudent55):

Good. Now remember that the derivative of a sum is the sum of the derivatives. \(\dfrac{d}{dx} [f(x) + g(x)] = \dfrac{d}{dx} f(x) + \dfrac{d}{dx} g(x) \)

OpenStudy (mathstudent55):

That means: \( \dfrac{dw}{dt} = \dfrac{d}{dt}[ (2t^2+2)(tant)-t] \) \( = \dfrac{d}{dt}[ (2t^2+2)(tant)] - \dfrac{d}{dt} t \)

OpenStudy (anonymous):

FS'+SF' (2t^2+2)(sec^2(x))+tan(t)(4t) - 1

OpenStudy (mathstudent55):

Now you need the product rule on the left part, and a simple derivative on the right part.

OpenStudy (mathstudent55):

I see that you are ahead of me. Exactly. That's correct.

OpenStudy (anonymous):

excellent, thank you

OpenStudy (mathstudent55):

Actually, the excellent goes to you! Good job. You're welcome.